OpenStudy (anonymous):

Simplify the expression below. -5-8

OpenStudy (anonymous):

-5-8 =-13

Parth (parthkohli):

-5 - (+8) = -(8+5) = -13

OpenStudy (anonymous):

You are basically subtracting 8 from -5 and when you do this you simply add those numbers but keep the negative sign.